A 0.5 hp, 6-pole induction motor is excited by a 3 –phase, 60 Hz source. If the full-load speed is

1140 rpm, calculate the slip


1
Expert's answer
2021-08-09T08:05:34-0400

We know that,

"Ns=120*\frac{f}{(P)}"

"=120*\frac{60}{(6)}=1200"

"slip=\frac{(Ns-Nr)}{(Ns)}"

"=\frac{(1200-1140)}{(1200)}"

"=0.05"